OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that contains the points (4, -7) and (0, 5).

OpenStudy (anonymous):

put the points (x1,y1) and (x2,y2) in the following equation: (y-y1)/(y2-y1)=(x-x1)/(x2-x1) then write in the standard form

OpenStudy (anonymous):

@DeviousDill can u do it now?

OpenStudy (anonymous):

I think i can 5-(-7)/0-4 5+7/-4 12/-4 -3

OpenStudy (anonymous):

(y - -7) (x - 4) ------- = ------ 0_o (5 - -7) (0 - 4)

OpenStudy (anonymous):

I'm so confused right now. I haven't ran into these equations since 8th grade. I don't remember this stuff. -.-

OpenStudy (anonymous):

am i correct

OpenStudy (anonymous):

I don't know If you're correct, because I don't even know how to do it...

OpenStudy (anonymous):

sorry,let me say

OpenStudy (anonymous):

(y+7/12)=(x-4)/-4

OpenStudy (anonymous):

hey plz look my answer n tell me if i am correct or not Sorous..

OpenStudy (anonymous):

now we have to wtite it in the standard form -4y-28=12x-48 12x+4y=-48+28 12x+4y=-20

OpenStudy (anonymous):

It only says that i have to write the equation in slope intercept form that contains those points

OpenStudy (anonymous):

the final answer: y=-3x-5

OpenStudy (anonymous):

slope=-3 intercept=-5

OpenStudy (anonymous):

@DeviousDill did u get it?

OpenStudy (anonymous):

@keskerem sorry,did u see the solution?

OpenStudy (anonymous):

I'm just so confused right now.

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

u have to points and u want to find the equation of the line so u use that formula

OpenStudy (anonymous):

Just about this whole thing. Solpe-intercept Equations....like how to find this answer. then the next question is... Write the equation in slope-intercept form of the line that has a slope of -2 and y-intercept of 2. I just don't get it

OpenStudy (anonymous):

y=ax+b a:slope b:y-intercept

OpenStudy (anonymous):

for an example u have slpoe( a=-2) and y-intercept(b=2)

OpenStudy (anonymous):

y = -2x + 2 ?

OpenStudy (anonymous):

so tellme wat is the line equation?

OpenStudy (anonymous):

good

OpenStudy (anonymous):

u r right

OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 3 and contains the point (2, 5). so... 5 = -2(3) + 2 ?

OpenStudy (anonymous):

no no

OpenStudy (anonymous):

*facepalm*

OpenStudy (anonymous):

wat does it mean?

OpenStudy (anonymous):

Slap

OpenStudy (anonymous):

uhum

OpenStudy (anonymous):

Why the facepalm @DeviousDill ???

OpenStudy (anonymous):

Cause i'm so confused. I get one, then I don't know how to do the next one. >_<

OpenStudy (anonymous):

@DeviousDill i'll help u,don worry

OpenStudy (anonymous):

for the next one u have to use this equation y-y1=a(x-x1) a=slope point=(x1,y1)

OpenStudy (anonymous):

I got the previous one, now I can't get this one... Write the equation in slope-intercept form of the line that has a slope of 3 and contains the point (2, 5).

OpenStudy (anonymous):

if your data is 2 point,u will use the first equation if ur data is a slope and a point u will use the last formula

OpenStudy (anonymous):

okay so, y-5 = 3(x - 2)

OpenStudy (anonymous):

good

OpenStudy (anonymous):

so write in this form:y=ax+b

OpenStudy (anonymous):

can u write?

OpenStudy (anonymous):

no =/

OpenStudy (anonymous):

y-5=3x-3*2 y=3x-6+5 y=3x-1

OpenStudy (anonymous):

@DeviousDill ,is it clear?

OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 3 and contains the point (2, 5). y-y1=a(x-x1) y - 5 = 3(x - 2) y - 5 = 3x - 6 +5 +5 y = 3x - 6 + 5 y = 3x -1 like that @sorour ?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so y = 3x - 1 is the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay this next one is weird.

OpenStudy (anonymous):

Using the slope-intercept method to graph this equation, in what order should these points on the line be drawn?  6(2x - 4y) = 3(2x + 4)